OpenStudy (anonymous):

The function for the cost of materials to make a shirt is f(x) = x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts. (1 point)

OpenStudy (anonymous):

So we know: \[\eqalign{ &f(x)=x+5 \\ &g(x)=5x+6 \\ }\] And they want us to find \(g[f(18)]\) So let us derive a function, \(h(x)=g[f(x)]\) \[h(x)=g[f(x)]=g(x+5)=5(x+5)+6=5x+25+6=5x+31\] Now the function evaluated at 18 is: \[h(18)=5(18)+31=90+31=121\]

OpenStudy (anonymous):

To check this, you can find \(f(18)\) and plug that value in for \(g(x)\): \[f(18)=18+5=23\] \[g(23)=5(23)+6=115+6=121\]

OpenStudy (anonymous):

Wait.. Wouldn't it be 30, not 31?

OpenStudy (anonymous):

Naw, \(25+6=31\)

OpenStudy (anonymous):

You're literally the best. I've been stuck on this forever.

OpenStudy (anonymous):

Haha I try Juli! Anytime :)